OpenStudy (anonymous):

Two events A nd B in a sample space such tht p(A)=0.5,p(B)=0.25 nd p(AUB) (a) r they mutually exclusive?why? (b) find p(A intersection (Bcomplement))?

OpenStudy (anonymous):

The question is missing the given value of p(AUB). In general: p(AUB) = p(A) + p(B) - p(A intersection B) Mutually exclusive if: p(AUB) = p(A) + p(B). That is, if p(A intersection B) = 0, which in English means: there is no chance that both A and B occur at the same time. This makes them mutually exclusive

OpenStudy (anonymous):

I m also confused

OpenStudy (anonymous):

and in general: p(A intersection B) = p(A) * p(B) and Bcompliment = 1 - B

OpenStudy (anonymous):

Plz help me i m havin exam :(

OpenStudy (anonymous):

i can't because your question is not complete. you seem to have omitted the end (before (a) and (b))

OpenStudy (anonymous):

can't answer (a) without knowing p(AUB)

OpenStudy (anonymous):

P(AUB) is givem

OpenStudy (anonymous):

to you but not to me. you forgot to write it

OpenStudy (anonymous):

Oh sorry

OpenStudy (anonymous):

P(AUB) is 0.70

OpenStudy (anonymous):

So I will copy/paste part of what I wrote above. Mutually exclusive if p(AUB) = p(A) + p(B) does p(AUB) = p(A) + p(B)?

OpenStudy (anonymous):

0.7 = 0.5 + 0.25 ?

OpenStudy (anonymous):

No becz its 0.75

OpenStudy (anonymous):

correct. so (a) is no

OpenStudy (anonymous):

Ok nd b??

OpenStudy (anonymous):

p(Bcompliment) = 1 - p(B) = 1 - 0.25 = 0.75 p(A intersection Bcomp) = p(A) * p(Bcomp) = 0.5*0.75 = 0.375

OpenStudy (anonymous):

Thank u @ euler271 :)

OpenStudy (anonymous):

@Euler271

OpenStudy (anonymous):

happy to help :)